# PPTA Collaboration

### v1

@@ IV PPA results and discussions
In Fig. 2, we also present two categories of additional limits as a reference. The first category does not rely on the ALDM assumption, including the limits from CAST Anastassopoulos et al. (2017), SN1987A Payez et al. (2015) and H1821+643 Reynés et al. (2021). Since the signals arise from relativistic axion-like particles, the generated limits are flat up to the energy scale of their production. The second category is to search for the ALDM-induced oscillations in polarization data, including the analyses of PPTA-QUIJOTE Castillo et al. (2022), CMB Ferguson et al. (2022), and Crab Nebula Adachi et al. (2024). The method in Ref. Fedderke et al. (2019) is based on washout effects caused by ALDM-induced oscillations on the CMB during the recombination era, which benefits from the relatively high DM density in the early universe. As these studies are ALDM-based, the derived limits on $g_{a\gamma\gamma}$ exhibit similar scaling features with $m_{a}$ to that of the PPPA limits. The PPPA provides the most stringent limits across a wide mass range within this category, and varying the value of $\rho_{0}$ does not change their relative positions. For a local value of $\rho_{0}=0.4\,\textrm{GeV}/\textrm{cm}^{3}$, especially, the PPPA offers the best global limits, namely $\lesssim 10^{-13.5}-10^{-12.2}~{}{\rm GeV}^{-1}$, for the mass range of fuzzy DM of $10^{-22}-10^{-21}~{}{\rm eV}$. Thanks to the 18-year observation time span of PPTA DR3, its excellent sensitivities extend to $\sim 10^{-23}\,$eV.
@@ V Conclusion and outlook
In this letter, we performed the first-ever PPA analysis to detect the ultralight ALDM, using the polarization data from the PPTA DR3. We developed a systematic analysis framework for this task, including proper data construction, advanced noise modeling and robust Bayesian sensitivity analysis. The derived limits are found to be superior to the existing ones for the mass range of fuzzy DM and hence represents the globally best. Particularly, the data correlation across the pulsars plays a pivotal role in recognizing the nature of these limits. This study represents an initial effort to investigate the great physical potential of current and forthcoming polarization data of pulsars. Besides further improvements of the developed PPA analysis framework, especially its noise modeling, several significant scientific tasks can be readily recognized for future explorations.
First of all, we can further advance the PPA applications for the ALDM detection. The PPA as an innovative tool is universal and can be built up with polarization data from other PTA programs also and even with broad datasets of normal pulsars. Moreover, we can synergize the PPA and the PTA. The ALDM halo can periodically perturb the spacetime metric and hence be detected by the PTA Khmelnitsky and Rubakov (2014); De Martino et al. (2017); Porayko et al. (2018); Afzal et al. (2023); Smarra et al. (2023). By constructing a PPA-PTA detector where the polarization and timing signals are correlated, the capability of a pulsar array could be further enhanced. Finally, the current PTA and PPA analyses could be extended from the radio band to high frequencies. With the pulsar data collected by Fermi Large Area Telescope, the first $\gamma$-ray PTA constraints on the SGWB have been delivered in Ajello et al. (2022) and on the ultralight ALDM in Luu et al. (2024); Xia et al. (2023). We would expect the current and forthcoming $X$-ray missions, such as the Imaging X-ray Polarimetry Explorer Weisskopf et al. (2022) and the Enhanced X-ray Timing and Polarimetry Zhang et al. (2019), to offer valuable insights on this task for the PPA. We leave these explorations to future work.
@@ Acknowledgments
X.X. is supported by Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC2121 “Quantum Universe” - 390833306. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya. X.X. was partly funded by the grant CNS2023-143767. Grant CNS2023-143767 funded by MICIU/AEI/10.13039/501100011033 and by European Union NextGenerationEU/PRTR. T.L. is supported by the Collaborative Research Fund under Grant No. C6017-20G which is issued by Research Grants Council of Hong Kong S.A.R. J.R. is supported in part by the National Natural Science Foundation of China under Grant No. 12275276. J.S. is supported by Peking University under startup Grant No. 7101302974 and the National Natural Science Foundation of China under Grants No. 12025507, No.12150015; and is supported by the Key Research Program of Frontier Science of the Chinese Academy of Sciences (CAS) under Grants No. ZDBS-LY-7003. Y.Z. is supported by U.S. Department of Energy under Award No. DESC0009959. Z.C.C. is supported by the National Natural Science Foundation of China (Grant No. 12247176 and No. 12247112) and the innovative research group of Hunan Province under Grant No. 2024JJ1006. Parts of this work were undertaken with the support of the Australian Research Council Centre of Excellence for Gravitational Wave Discovery (CE230100016). L.Z. is supported by ACAMAR Postdoctoral Fellowship and the National Natural Science Foundation of China (Grant No. 12103069). XJZ is supported by the National Natural Science Foundation of China (Grant No. 12203004) and by the Fundamental Research Funds for the Central Universities.
@@ VI Supplemental Materials
The original data of PPTA DR3 contains the measurements of four Stokes parameters in pulsar fold-modes, spanning approximately from 2004 to 2022 Zic et al. (2023). The data from Ultra-Wideband Low (UWL) receiver Hobbs et al. (2020) are accumulated since 2018. We reprocess the UWL data set using the most up-to-date polarization model following procedures described in Hobbs et al. (2020). The Stokes parameters are measured as an average over either 0.5 or $1$ hour at each epoch. After subtracting Faraday rotation caused by the ISM, we average the observation for each epoch over the considered frequency band. We finally have Stokes parameters as a function of the pulse phase and observational epoch. Note that, for the ALDM mass of interest, i.e., $m_{a}\lesssim 10^{-18}\,$eV, the induced signal oscillates with a period longer than hours. The relevant information is thus expected to be well-preserved by the data processed in this procedure.
@@
In this article, we focus our analysis on the frequency band at $10$cm, which has a central frequency of $3100$ MHz with a bandwidth of 1GHz. This frequency band is relatively high for the PPTA data, subjecting to relatively weak Faraday rotation. To keep this choice consistent, we use the data only within the same frequency range for the UWL dataset. Finally, we chose 22 MSPs out of 32 in PPTA DR3 to construct the PPPA with significant linear polarization SNR in this frequency band. The observations of PSR J1713+0747 during its recent event of pulse profile change Xu et al. (2021); Singha et al. (2021); Jennings et al. (2024) have been excluded. We display the basic information on these PPPA pulsars in Table 1.
Pulsar Observations Pulsar Distance med SD Refs
$T_{\rm obs}$(year) $N_{\rm obs}$ ${\rm max}(\tilde{L})/\sigma_{\tilde{L}}$ $\widetilde{\rm DM}$ (kpc) PX (mas) ($\sigma_{n}/$deg) ($\Delta{\rm PA}/$deg)
1 J0437-4715 18.0 1122 976.4 0.156 6.37(9) 0.13 0.91 Johnston et al. (1993); Reardon et al. (2016); Yao et al. (2017a)
2 J0613-0200 17.5 333 17.9 1.024 1.01(9) 8.0 8.2 Lorimer et al. (1995); Reardon et al. (2021); Yao et al. (2017a)
3 J0614-3329 2.8 37 42.8 2.691 1.1(11) 4.5 3.2 Ransom et al. (2010); Curyło et al. (2023); Yao et al. (2017a)
4 J0711-6830 17.8 371 67.2 0.106 N/A 7.7 9.0 Bailes et al. (1997); Reardon et al. (2021); Yao et al. (2017a)
5 J1017-7156 11.4 348 112.0 1.807 0.6(6) 3.5 4.1 Keith et al. (2012); Reardon et al. (2021); Yao et al. (2017a)
6 J1022+1001 17.9 334 579.0 0.834 1.55(16) 0.86 1.2 Camilo et al. (1996); Reardon et al. (2021); Yao et al. (2017a)
7 J1024-0719 17.8 221 142.8 0.382 0.93(5) 1.9 1.9 Bailes et al. (1997); Reardon et al. (2021); Yao et al. (2017a)
8 J1045-4509 16.7 274 77.0 0.334 1.7(7) 2.5 2.6 Bailes et al. (1994); Reardon et al. (2021); Yao et al. (2017a)
9 J1125-6014 7.4 164 187.7 0.988 0.6(33) 2.7 2.2 Faulkner et al. (2004); Reardon et al. (2021); Yao et al. (2017a)
10 J1545-4550 10.1 185 438.7 2.247 0.45(14) 0.92 1.1 Burgay et al. (2013); Reardon et al. (2021); Yao et al. (2017a)
11 J1600-3053 18.0 305 355.9 2.535 0.53(6) 0.98 1.3 Jacoby et al. (2007); Reardon et al. (2021); Yao et al. (2017a)
12 J1603-7202 18.0 293 110.7 1.129 0.3(3) 4.7 5.4 Lorimer et al. (1996); Reardon et al. (2021); Yao et al. (2017a)
13 J1643-1224 17.9 274 128.1 0.791 1.1(1) 1.9 2.3 Lorimer et al. (1995); Reardon et al. (2021); Yao et al. (2017a)
14 J1713+0747 17.1 319 1400.6 0.919 0.763(21) 0.43 1.2 Foster et al. (1993); Reardon et al. (2021); Yao et al. (2017a)
15 J1730-2304 17.9 226 530.9 0.512 2.0(1) 0.86 1.4 Lorimer et al. (1995); Reardon et al. (2021); Yao et al. (2017a)
16 J1744-1134 18.0 329 1206.9 0.148 2.44(5) 0.39 0.72 Bailes et al. (1997); Reardon et al. (2021); Yao et al. (2017a)
17 J1824-2452A 15.8 120 241.7 3.702 0.200(11) 1.1 1.1 Lyne et al. (1987); Reardon et al. (2021); Yao et al. (2017a)
18 J1857+0943 17.6 182 71.8 0.767 0.85(16) 6.0 6.5 Segelstein et al. (1986); Reardon et al. (2021); Yao et al. (2017a)
19 J1909-3744 18.0 918 2078.6 0.564 0.86(1) 0.71 1.8 Jacoby et al. (2003); Reardon et al. (2021); Yao et al. (2017a)
20 J1939+2134 16.3 162 77.4 2.896 0.35(3) 2.0 1.9 Backer et al. (1982); Reardon et al. (2021); Yao et al. (2017a)
21 J2145-0750 17.8 299 140.6 0.693 1.40(8) 2.9 3.1 Bailes et al. (1994); Reardon et al. (2021); Yao et al. (2017a)
22 J2241-5236 12.0 355 119.9 0.963 0.96(4) 5.2 4.6 Keith et al. (2011); Reardon et al. (2021); Yao et al. (2017a)
@@ VII.1 Time series of the PA residuals
[…]
@@
Figure 7 displays the time series of PA residuals for the 22 PPPA MSPs. In the analysis, we discard a small number of observations where the PA residuals deviate by more than ±$\pm 4\sigma$ from the baseline or have an error bar larger than $4\sigma$ or $30$ degrees. These observations account for less than 5% of total observations for all pulsars.
@@
One major contribution to the observed PA residuals is Faraday rotation induced by the Earth’s ionosphere. This effect can be described as $\Delta{\rm PA}^{\rm ion}=\phi^{\rm ion}\lambda^{2}$, and is important for radio band. Here $\phi^{\rm ion}$ is ionospheric rotation measure (RM). For a long time, the ionospheric effects have been calculated with the numerical code FARROT, developed at the Dominion Radio Astrophysical Observatory, and the International Reference Ionosphere (IRI) model Han et al. (2006). We instead employ ionFR Sotomayor-Beltran et al. (2013), a program developed more recently to calculate the ionospheric RM using the International Geomagnetic Reference Field (IGRF-13) model and publicly available global electron maps. We provide the coordinates and observation epochs of the PPPA pulsars and the Earth’s position of Parkes observatory as inputs to ionFR to determine the line-of-sight to each observation and compute its ionospheric RM. As the ionospheric data is available on an hourly basis, a binning approach is applied for this calculation. Concretely, if the median time of an observation epoch is between 30 minutes before and after an exact hour, we subtract the observed PA residual by the ionospheric RM of this hour. This hourly timescale sets a cutoff for the sensitivity range at the large mass end.
@@
The ionospheric Faraday rotation is strongly influenced by solar activities, and varies significantly over time, on scales of hours to years, and direction. These effects complicate the prediction and yield a large uncertainty for the predicted RM values. For instance, the ionFR prediction for a given pulsar may fluctuate by up to $\sim 1$ rad m-2 within two hours, particularly during intense solar activities, and can also differ from the IRI one by $\mathcal{O}(1)$ rad m-2. Furthermore, for many pulsars, we observe a discrepancy between the median measurement error and the SD of PA residuals after ionospheric subtraction (see the last two columns in Table 1). This suggests the presence of unmodeled noise in the PA residual time series, despite the efforts to model ionospheric corrections using ionFR. To account for these observations, we incorporate additional random and deterministic noise components alongside radiometer noise and modeled ionospheric corrections, following the common practices in the PTA analysis Lentati et al. (2014). This advanced noise modeling has not been considered in the existing analyses of pulsar polarization data, and turns out to be crucial for optimizing the PPA detection of the ALDM signal.
@@
Here $\sigma_{p,n}$ is the measurement error of the observed PA residual $\Delta{\rm PA}^{\rm obs}_{p,n}$ for the $n$-th epoch of the $p$-th pulsar. EFp is a rescaling factor introduced to model the non-Gaussianity of the observational noise, and EQp is additional white noise added in quadrature. We treat these two factors as free for individual pulsars. Their prior distributions are given by
@@
where ${\rm U}$ is a uniform distribution. Note, EQp has been commonly used for modeling jitter noise, namely random pulse-to-pulse variation, in high-precision pulsar timing analysis Shannon et al. (2014). For the pulsar polarization data analyzed here, the jitter noise is not significant as a contribution to the observed PA residuals. Actually, the measured jitter noise in the considered pulsars yields a profile variation by only up to $\sim 1/100$ of a phase bin. Instead, a major source for additional white noise may arise from ionospheric Faraday rotation on timescales much shorter than the observation cadence.
@@
with a spacing of $\Delta f=1/T_{{\rm obs},p}$. While the approach to address red noise here is similar to that used in the PTA studies Agazie et al. (2023b), the dominant source may differ. One potential source is the ISM whose RM vary linearly over time with a slope of $\sim 0.1-1$ rad m-2 per year, correlated with $\widetilde{\rm DM}$ Keith et al. (2024). Given relatively small RM and $\widetilde{\rm DM}$ magnitudes for many PPPA pulsars, the ISM contribution to red noise might be minimal. The ionospheric Faraday rotation may vary on a timescale of years and thus contribute more to red noise for these pulsars.
For additional deterministic contributions, we observe temporal variations in data at a long time scale, in addition to a constant offset. This effect could be related to slow changes in ISM properties Keith et al. (2024). Without loss of generality, we model the deterministic PA residuals as a polynomial up to the second order, namely
$\Delta{\rm PA}^{\rm det}_{p}(t)=\psi^{(0)}_{p}+\psi^{(1)}_{p}t+\psi^{(2)}_{p}t^{2}\,,$ (25)
where $\psi^{(0)}_{p},\psi^{(1)}_{p},\psi^{(2)}_{p}$ are the coefficients of intrinsic, linear and quadratic terms, respectively. For a given pulsar, $\Delta{\rm PA}^{\rm det}_{p}(t)$ can be recast as a vector of deterministic PA residuals in terms of the relevant observation epochs $\{t_{p,n}\}$:
$\Delta\boldsymbol{{\rm PA}}^{\rm det}_{p}=\boldsymbol{M}_{p}\boldsymbol{\psi}_{p}$ (26)
@@
$\Delta\boldsymbol{{\rm PA}}^{\rm det}$ $=$ $\begin{pmatrix}\boldsymbol{{\rm PA}}^{\rm det}_{1}\\ \boldsymbol{{\rm PA}}^{\rm det}_{2}\\ \vdots\\ \boldsymbol{{\rm PA}}^{\rm det}_{\mathcal{N}}\end{pmatrix},\quad\boldsymbol{\psi}=\begin{pmatrix}\boldsymbol{\psi}_{1}\\ \boldsymbol{\psi}_{2}\\ \vdots\\ \boldsymbol{\psi}_{\mathcal{N}}\end{pmatrix},$

### v2

@@ IV PPA results and discussions
In Fig. 2, we also present two categories of additional limits as a reference. The first category does not rely on the ALDM assumption, including the limits from CAST Anastassopoulos et al. (2017), SN1987A Payez et al. (2015) and H1821+643 Reynés et al. (2021). Since the signals arise from relativistic axion-like particles, the generated limits are flat up to the energy scale of their production. The second category is to search for the ALDM-induced oscillations in polarization data, including the analyses of PPTA-QUIJOTE Castillo et al. (2022), CMB Ferguson et al. (2022), and Crab Nebula Adachi et al. (2024). The method in Ref. Fedderke et al. (2019) is based on washout effects caused by ALDM-induced oscillations on the CMB during the recombination era, which benefits from the relatively high DM density in the early universe. As these studies are ALDM-based, the derived limits on $g_{a\gamma\gamma}$ exhibit similar scaling features with $m_{a}$ to that of the PPPA limits. The PPPA provides the most stringent limits across a wide mass range within this category, and varying the value of $\rho_{0}$ does not change their relative positions. For a local density value of $\rho_{0}=0.4\,\textrm{GeV}/\textrm{cm}^{3}$, especially, the PPPA offers the best global limits, namely $\lesssim 10^{-13.5}-10^{-12.2}~{\rm GeV}^{-1}$, for the mass range of fuzzy DM of $10^{-22}-10^{-21}~{\rm eV}$. Thanks to the 18-year observation time span of PPTA DR3, its excellent sensitivities extend to $\sim 10^{-23}\,$eV.
@@ V Conclusion and outlook
In this Letter, we presented the first-ever PPA analysis to detect the ultralight ALDM, using the polarization data from the PPTA DR3. We developed a systematic analysis framework for this task, including proper data construction, advanced noise modeling and robust Bayesian sensitivity analysis. The derived limits are found to be superior to the existing ones for the mass range of fuzzy DM and hence represents the globally best. Particularly, the data correlation across the pulsars plays a pivotal role in recognizing the nature of these limits. This study represents an initial effort to investigate the great physical potential of current and forthcoming polarization data of pulsars. Besides further improvements of the developed PPA analysis framework, especially its noise modeling, several significant scientific tasks can be readily recognized for future explorations.
First of all, we can further advance the PPA applications for the ALDM detection. The PPA as an innovative tool is universal and can be built up with polarization data from other PTA programs also and even with broad datasets of normal pulsars. Moreover, we can synergize the PPA and the PTA. The ALDM halo can periodically perturb the spacetime metric and hence be detected by the PTA Khmelnitsky and Rubakov (2014); De Martino et al. (2017); Porayko et al. (2018); Afzal et al. (2023); Smarra et al. (2023). By constructing a PPA-PTA detector where the polarization and timing signals are correlated, the capability of a pulsar array could be further enhanced Li et al. (2025). Finally, the current PTA and PPA analyses could be extended from the radio band to high frequencies. With the pulsar data collected by Fermi Large Area Telescope, the first $\gamma$-ray PTA constraints on the SGWB have been delivered in Ajello et al. (2022) and on the ultralight ALDM in Luu et al. (2024); Xia et al. (2023). We would expect the current and forthcoming x-ray missions, such as the Imaging X-ray Polarimetry Explorer Weisskopf et al. (2022) and the Enhanced X-ray Timing and Polarimetry Zhang et al. (2019), to offer valuable insights on this task for the PPA. We leave these explorations to future work.
@@ Acknowledgments
X.X. is supported by Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC2121 “Quantum Universe” - 390833306. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya. X.X. was partly funded by the grant CNS2023-143767. Grant CNS2023-143767 funded by MICIU/AEI/10.13039/501100011033 and by European Union NextGenerationEU/PRTR. T.L. is supported by the Collaborative Research Fund under Grant No. C6017-20G which is issued by Research Grants Council of Hong Kong S.A.R. J.R. is supported in part by the National Natural Science Foundation of China under Grant No. 12435005. J.S. is supported by Peking University under startup Grant No. 7101302974 and the National Natural Science Foundation of China under Grants No. 12025507, No.12150015; and is supported by the Key Research Program of Frontier Science of the Chinese Academy of Sciences (CAS) under Grants No. ZDBS-LY-7003. Y.Z. is supported by U.S. Department of Energy under Award No. DESC0009959. Z.C.C. is supported by the National Natural Science Foundation of China (Grant No. 12247176 and No. 12247112) and the innovative research group of Hunan Province under Grant No. 2024JJ1006. Parts of this work were undertaken with the support of the Australian Research Council Centre of Excellence for Gravitational Wave Discovery (CE230100016). L.Z. is supported by ACAMAR Postdoctoral Fellowship and the National Natural Science Foundation of China (Grant No. 12103069). XJZ is supported by the National Natural Science Foundation of China (Grant No. 12203004) and by the Fundamental Research Funds for the Central Universities.
@@ VI Supplemental Materials
The original data of PPTA DR3 contains the measurements of four Stokes parameters in pulsar fold-modes, spanning approximately from 2004 to 2022 Zic et al. (2023). The data from Ultra-Wideband Low (UWL) receiver Hobbs et al. (2020) are accumulated since 2018. We reprocess the UWL data set using the most up-to-date polarization model following procedures described in Hobbs et al. (2020). Radio frequency interference (RFI) has been identified, masked, and flagged as part of the PPTA DR3 process in Zic et al. (2023), whose data products were utilized in this study. A two-stage RFI mitigation procedure was applied to the UWL data sets to reduce interference across the wide frequency band, while the data sets prior to UWL were reprocessed using more advanced RFI excision techniques Zic et al. (2023). [^1] The RFI is complex due to a mixture of diverse sources, including radar, satellites, and various local electronic facilities. While a mitigation procedure was applied to remove easily identifiable RFI from the data, we anticipate that more subtle RFI may still be present in the datasets, contributing to both temporally uncorrelated and correlated noise. We leave this investigation to a future work.
@@
The Stokes parameters are measured as an average over either 0.5 or $1$ hour at each epoch. After subtracting Faraday rotation caused by the ISM, we average the observation for each epoch over the considered frequency band. We finally have Stokes parameters as a function of the pulse phase and observational epoch. Note that, for the ALDM mass of interest, i.e., $m_{a}\lesssim 10^{-18}\,$eV, the induced signal oscillates with a period longer than hours. The relevant information is thus expected to be well-preserved by the data processed in this procedure.
@@
In this article, we focus our analysis on the frequency band at 1010cm, which has a central frequency of $3100$ MHz with a bandwidth of 1GHz. This frequency band is relatively high for the PPTA data, subjecting to relatively weak Faraday rotation. To keep this choice consistent, we use the data only within the same frequency range for the UWL dataset. Finally, we chose 22 MSPs out of 32 in PPTA DR3 to construct the PPPA with significant linear polarization SNR in this frequency band. The observations of PSR J1713+0747 during its recent event of pulse profile change Xu et al. (2021); Singha et al. (2021); Jennings et al. (2024) have been excluded. We display the basic information on these PPPA pulsars in Table 1.
Pulsar Observations Ecliptic Latitude Pulsar Distance med SD Refs
$T_{\rm obs}$(year) $N_{\rm obs}$ ${\rm max}(\tilde{L})/\sigma_{\tilde{L}}$ (deg) $\widetilde{\rm DM}$ (kpc) PX (mas) ($\sigma_{n}/$deg) ($\Delta{\rm PA}/$deg)
1 J0437–4715 18.0 1122 976.4 -67.87 0.156 6.37(9) 0.13 0.91 Johnston et al. (1993); Reardon et al. (2016); Yao et al. (2017a)
2 J0613–0200 17.5 333 17.9 -25.41 1.024 1.01(9) 8.0 8.2 Lorimer et al. (1995); Reardon et al. (2021); Yao et al. (2017a)
3 J0614–3329 2.8 37 42.8 -56.87 2.691 1.1(11) 4.5 3.2 Ransom et al. (2010); Curyło et al. (2023); Yao et al. (2017a)
4 J0711–6830 17.8 371 67.2 -82.89 0.106 N/A 7.7 9.0 Bailes et al. (1997); Reardon et al. (2021); Yao et al. (2017a)
5 J1017–7156 11.4 348 112.0 -67.74 1.807 0.6(6) 3.5 4.1 Keith et al. (2012); Reardon et al. (2021); Yao et al. (2017a)
6 J1022+1001 17.9 334 579.0 -0.06 0.834 1.55(16) 0.86 1.2 Camilo et al. (1996); Reardon et al. (2021); Yao et al. (2017a)
7 J1024–0719 17.8 221 142.8 -16.04 0.382 0.93(5) 1.9 1.9 Bailes et al. (1997); Reardon et al. (2021); Yao et al. (2017a)
8 J1045–4509 16.7 274 77.0 -47.71 0.334 1.7(7) 2.5 2.6 Bailes et al. (1994); Reardon et al. (2021); Yao et al. (2017a)
9 J1125–6014 7.4 164 187.7 -55.66 0.988 0.6(33) 2.7 2.2 Faulkner et al. (2004); Reardon et al. (2021); Yao et al. (2017a)
10 J1545–4550 10.1 185 438.7 -25.29 2.247 0.45(14) 0.92 1.1 Burgay et al. (2013); Reardon et al. (2021); Yao et al. (2017a)
11 J1600–3053 18.0 305 355.9 -10.07 2.535 0.53(6) 0.98 1.3 Jacoby et al. (2007); Reardon et al. (2021); Yao et al. (2017a)
12 J1603–7202 18.0 293 110.7 -49.96 1.129 0.3(3) 4.7 5.4 Lorimer et al. (1996); Reardon et al. (2021); Yao et al. (2017a)
13 J1643–1224 17.9 274 128.1 9.78 0.791 1.1(1) 1.9 2.3 Lorimer et al. (1995); Reardon et al. (2021); Yao et al. (2017a)
14 J1713+0747 17.1 319 1400.6 30.70 0.919 0.763(21) 0.43 1.2 Foster et al. (1993); Reardon et al. (2021); Yao et al. (2017a)
15 J1730–2304 17.9 226 530.9 0.19 0.512 2.0(1) 0.86 1.4 Lorimer et al. (1995); Reardon et al. (2021); Yao et al. (2017a)
16 J1744–1134 18.0 329 1206.9 11.81 0.148 2.44(5) 0.39 0.72 Bailes et al. (1997); Reardon et al. (2021); Yao et al. (2017a)
17 J1824–2452A 15.8 120 241.7 -1.55 3.702 0.200(11) 1.1 1.1 Lyne et al. (1987); Reardon et al. (2021); Yao et al. (2017a)
18 J1857+0943 17.6 182 71.8 32.32 0.767 0.85(16) 6.0 6.5 Segelstein et al. (1986); Reardon et al. (2021); Yao et al. (2017a)
19 J1909–3744 18.0 918 2078.6 -15.16 0.564 0.86(1) 0.71 1.8 Jacoby et al. (2003); Reardon et al. (2021); Yao et al. (2017a)
20 J1939+2134 16.3 162 77.4 42.30 2.896 0.35(3) 2.0 1.9 Backer et al. (1982); Reardon et al. (2021); Yao et al. (2017a)
21 J2145–0750 17.8 299 140.6 5.31 0.693 1.40(8) 2.9 3.1 Bailes et al. (1994); Reardon et al. (2021); Yao et al. (2017a)
22 J2241–5236 12.0 355 119.9 -40.39 0.963 0.96(4) 5.2 4.6 Keith et al. (2011); Reardon et al. (2021); Yao et al. (2017a)
@@ VII.1 Time series of the PA residuals
where $n_{n,k}$ is observational noise, with $\langle n_{n,k}\rangle=0$ (here and below $\langle\cdot\cdot\cdot\rangle$ denotes an average over an ensemble of possible realizations for a given pulsar), and
@@
Figure 8: Time series of the observed PA residuals in Eq. (13) for the 22 PPPA MSPs, after the ionospheric subtraction with ionFR. The central value of each observation and its error bar are denoted in blue and red, respectively.
Fig. 7 displays the time series of PA residuals for the 22 PPPA MSPs. In the analysis, we discard a small number of observations where the PA residuals deviate by more than $\pm 4\sigma$ from the baseline or have an error bar larger than $4\sigma$ or $30$ degrees. These observations account for less than 5% of total observations for all pulsars. One observation in Fig. 7 is the annual oscillations in $\Delta{\rm PA}^{\rm obs}$ for pulsars such as PSRs J1713+0747 and J1744-1134. These oscillations are likely driven by variations in ionospheric total-electron content Sotomayor-Beltran et al. (2013), rather than by solar elongation. Actually, one can see from Table 1 that PSRs J1713+0747 and J1744-1134 have ecliptic latitudes of 30.7 and 11.8, respectively, much higher than those of many other pulsars in our sample which do not show such a time-varying trend. In contrast, J1022+1001 has the lowest ecliptic latitude in our sample at $0.1^{\circ}$, yet showing no significant annual variation. A detailed discussion on the ionospheric effects and the noises that they may introduce on various timescales is provided in Sec. B.
@@
One major contribution to the observed PA residuals is Faraday rotation induced by the Earth’s ionosphere. This effect can be described as $\Delta{\rm PA}^{\rm ion}=\phi^{\rm ion}\lambda^{2}$, and is important for radio band. Here $\phi^{\rm ion}$ is ionospheric rotation measure (RM). For a long time, the ionospheric effects have been calculated with the numerical code FARROT, developed at the Dominion Radio Astrophysical Observatory, and the International Reference Ionosphere (IRI) model Han et al. (2006). We instead employ ionFR Sotomayor-Beltran et al. (2013), a program developed more recently to calculate the ionospheric RM using the International Geomagnetic Reference Field (IGRF-13) model and publicly available global electron maps. We provide the coordinates and observation epochs of the PPPA pulsars and the Earth’s position of Parkes observatory as inputs to ionFR to determine the line-of-sight to each observation and compute its ionospheric RM. As the ionospheric data is available on an hourly basis, a binning approach is applied for this calculation. Concretely, if the median time of an observation epoch is between 30 minutes before and after an exact hour, we subtract the observed PA residual by the ionospheric RM of this hour. This hourly timescale sets a cutoff for the sensitivity range at the large mass end. In Fig. 8, we demonstrate the time series of PA residuals for the 22 PPPA MSPs, after the ionospheric subtraction with ionFR.
@@
The ionospheric Faraday rotation is strongly influenced by solar activities, and its contributions to the observed PA residuals may vary significantly over time, on scales of hours to years, and direction. For example, solar activities can greatly impact the ionospheric total electron content, leading to considerable annual variations in the observed PA residuals. As illustrated in Fig. 8, the apparent annual oscillations for PSRs J1713+0747 and J1744-1134 in Fig. 7 get significantly reduced - though not entirely eliminated - supporting the conjecture of their ionospheric origin  [^2] The observed annual oscillations for these pulsars may receive contributions also from the ISM effects Hotan et al. (2004); Everett and Weisberg (2001); Arzoumanian et al. (2023), such as the annual variations of magnetic field strength caused by magnetized clumps of material along the line of sight to the pulsars Arzoumanian et al. (2023).. Generally, the complexity of solar activities make it challenging to predict the ionospheric effects and hence yield a large uncertainty for the predicted RM values. The ionFR prediction for a given pulsar may fluctuate by up to $\sim 1$ rad m-2 within two hours, particularly during intense solar activities, and can also differ from the IRI one by $\mathcal{O}(1)$ rad m-2. Furthermore, for many pulsars, we observe a discrepancy between the median measurement error and the SD of PA residuals after ionospheric subtraction (see the last two columns in Table 1). This suggests the presence of unmodeled noise in the PA residual time series, despite the efforts to model ionospheric corrections using ionFR. To account for these observations, we incorporate additional noise components alongside radiometer noise and modeled ionospheric corrections, following the common practices in the PTA analysis Lentati et al. (2014). This advanced noise modeling has not been considered in the existing analyses of pulsar polarization data, and turns out to be crucial for optimizing the PPA detection of the ALDM signal.
@@
Here $\sigma_{p,n}$ is the measurement error of the observed PA residual $\Delta{\rm PA}^{\rm obs}_{p,n}$ for the $n$-th epoch of the $p$-th pulsar, EFp is a rescaling factor of the measurement error, and EQp is additional white noise added in quadrature. EFp and EQp are free parameters. We introduce them to collectively account for the contributions from uncorrelated and weakly correlated processes, considering the complexity of measuring the individual components of white noise. Their prior distributions are given by
@@
where $\rm U$} is a uniform distribution. Note, EQp has been commonly used for modeling jitter noise, namely random pulse-to-pulse variation, in high-precision pulsar timing analysis Shannon et al. (2014). For the pulsar polarization data analyzed here, the jitter noise is not significant as a contribution to the observed PA residuals. Actually, the measured jitter noise in the considered pulsars yields a profile variation by only up to $\sim 1/100$ of a phase bin. However, complex effects such as underestimated or overestimated error bars, $\Delta{\rm PA}$ variations at short timescales due to unmodeled ionospheric effects, environmental fluctuations, and even residual RFI effects after the mitigation procedure Zic et al. (2023) could bring in additional white noise and contribute to EFp and EQp.
@@
with a spacing of $\Delta f=1/T_{{\rm obs},p}$. The approach to address red noise here is similar to that used for the PTA studies Agazie et al. (2023b). The major source could be the ISM where the RM varies linearly over time with a slope $\sim 0.1-1$ rad m-2 per year, correlated with $\widetilde{\rm DM}$ Keith et al. (2024). Additionally, unmodeled ionospheric effects after the ionFR subtraction may also vary over time and contribute to the red noise for some pulsars.
To account for potential DM and RM variations of long timescale, we incorporate a polynomial component of up to second-order in the noise model:
$\Delta{\rm PA}^{\rm pol}_{p}(t)=\psi^{(0)}_{p}+\psi^{(1)}_{p}t+\psi^{(2)}_{p}t^{2}\,$, (25)
where $\psi^{(0)}_{p},\psi^{(1)}_{p},\psi^{(2)}_{p}$ are the coefficients of intrinsic, linear and quadratic terms, respectively. These temporal variations could arise from slow changes in ISM properties, particularly due to gradual stochastic variations in electron density included in the DM and RM Lam et al. (2016); Keith et al. (2024). For a given pulsar, $\Delta{\rm PA}^{\rm pol}_{p}(t)$ can be recast as a vector of its values at the relevant observation epochs $\{t_{p,n}\}$:
$\Delta\boldsymbol{{\rm PA}}^{\rm pol}_{p}=\boldsymbol{M}_{p}\boldsymbol{\psi}_{p}$ (26)
@@
$\Delta\boldsymbol{{\rm PA}}^{\rm pol}$ $=$ $\begin{pmatrix}\Delta\boldsymbol{{\rm PA}}^{\rm pol}_{1}\\ \Delta\boldsymbol{{\rm PA}}^{\rm pol}_{2}\\ \vdots\\ \Delta\boldsymbol{{\rm PA}}^{\rm pol}_{\mathcal{N}}\end{pmatrix},\quad\boldsymbol{\psi}=\begin{pmatrix}\boldsymbol{\psi}_{1}\\ \boldsymbol{\psi}_{2}\\ \vdots\\ \boldsymbol{\psi}_{\mathcal{N}}\end{pmatrix}$,